OpenStudy (anonymous):

What is the x-intercept in this linear equation?2y - 4x = -12

OpenStudy (luigi0210):

set y equal to zero bro,and solve

OpenStudy (anonymous):

2y-4x=-12 2(0)-4x=-12 4x=-12 x=-3

OpenStudy (luigi0210):

Ah, guess what? you lost your negative there

OpenStudy (anonymous):

do i take out the negative

OpenStudy (luigi0210):

Nope, keep them. 2(0)-4x=-12 -4x=-12 There now finish it